Circular No. 6699

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cw IN NGC 105 Q.-y. Qiao, Y.-l. Qiu, W.-d. Li, W. Zhou and J.-y. Hu, on behalf of the

Beijing Astronomical Observatory (BAO) Supernova Survey, report the discovery of a supernova of mag ∼ 16.5 on an unfiltered CCD frame taken on July 10.79 UT with the BAO 0.60-m reflector. A confirming image obtained on July 11.70 shows the object at mag ∼ 16.4. SN 1997cw is located at α = 0^{h}25^{m}17^{s}.27, δ = +12^{o}53′06′′.2 (equinox 2000.0), which is 7′′.6 east and 4′′.2 north of the nucleus of NGC 105. An unfiltered CCD frame taken on 1996 Nov. 3 (limiting mag ∼ 19.0) shows no star at this position. A low-resolution spectrum taken on July 11.8 by Qiu, Qiao, Li, Hu and J.-y. Wei indicates that SN 1997cw is an SN 1991T-like peculiar type Ia SN, near or soon after optical maximum. The entire spectrum resembles that of SN 1997br some three days after maximum. Weak, yet conspicuous Si II 635.5-nm absorption is present at 622 nm. Strong Na D absorption (equivalent width 3.7 nm) from the host galaxy suggests that the supernova is seriously reddened by NGC 105. There is also conspicuous Hα emission from the host galaxy, for which they estimate the redshift to be ∼ 4600 km/s.

GEMINGA H. O. Vats, M. R. Deshpande, C. Shah, A. K. Singal, K. N. Iyer,

R. Oza and S. Doshi, Physical Reasearch Laboratory, Ahmedabad, report the detection of radio pulses on June 19 from Geminga = 1E 0630+178 at a period of 237 ms, with a subpulse at 0.5 phase separation. The observations were made at 103 MHz using the 5000 sq. meter radio telescope at Rajkot.

CTA 102 M. T. Carini and J. C. Noble, Western Kentucky University; H. R.

Miller, Georgia State University; and T. J. Balonek, E. Beckerman and S. M. Davis, Foggy Bottom Observatory (FBO), Colgate University, report: “We have obtained R-band observations of CTA 102 (cf. IAUC 5477) with the 1.1-m Hall (H) telescope at the Lowell Observatory and the 0.4-m telescope at FBO, in response to reports of a flare at millimeter wavelengths: May 18, 16.31 ± 0.05 (FBO); 19, 16.07 ± 0.10 (H); 26, 15.98 ± 0.05 (FBO); 27, 16.02 ± 0.03 (H); 28, 16.17 ± 0.05 (H); 29, 16.11 ± 0.02 (H); June 4, 15.54 ± 0.02 (FBO); 10, 15.29 ± 0.03 (FBO); 11, 15.51 ± 0.02 (FBO); 16, 15.59 ± 0.03 (FBO).”

1997 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Timothy B. Spahr